# Generalized Moving Least-Squares for Solving Vector-valued PDEs on Unknown Manifolds

### v1

@@ B.2 L Laplacian
We first review the extrinsic formulation for the L Laplacian (see e.g. [52, 22]). Let $v={\mathrm{grad}}_{g}u$ be the gradient of a vector field $u$ and let $V$ be a smooth extension of $v$ onto an open subset in $\mathbb{R}^{3}$ such that $V|_{M}=v$. Based on the definition of the L Laplacian in (44), one can obtain the extrinsic formulation for the L Laplacian:
$\Delta_{L}u$ $=$ $\mathrm{div}_{g}\left({v+v}^{\top}\right)=\mathbf{P}\mathrm{tr}^{2}_{1}\left(\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}(V+V^{\top})\right)=\mathbf{P}\left[\begin{array}[]{c}\sum_{k=1}^{3}\mathcal{G}_{k}\left(V^{1k}+V^{k1}\right)\\ \sum_{k=1}^{3}\mathcal{G}_{k}\left(V^{2k}+V^{k2}\right)\\ \sum_{k=1}^{3}\mathcal{G}_{k}\left(V^{3k}+V^{k3}\right)\end{array}\right]$ (76)
$=$ $\mathcal{H}_{1}\left[V\mathbf{|}_{1}+\left(\bar{V}_{1}\right)^{\top}\right]+\mathcal{H}_{2}\left[V\mathbf{|}_{2}+\left(\bar{V}_{2}\right)^{\top}\right]+\mathcal{H}_{3}\left[V\mathbf{|}_{3}+\left(\bar{V}_{3}\right)^{\top}\right],$ (77)
where $\bar{V}_{i}=\left[V_{i1},V_{i2},V_{i3}\right]$ is the $i$th row of the $(2,0)$ tensor field $V$ for $i=1,2,3$. Here, we have used the notations $\mathcal{G}_{j}$ in (22) and $\mathcal{H}_{j}$ in (25) for $j=1,2,3$, and we can further calculate $\bar{V}^{i}$ as
$(\bar{V}^{i})^{\top}$ $=$ $\left[V_{i1},V_{i2},V_{i3}\right]^{\top}=\left(\left[P_{i1},P_{i2},P_{i3}\right]\left[\begin{matrix}\mathcal{G}_{1}U^{1}&\mathcal{G}_{2}U^{1}&\mathcal{G}_{3}U^{1}\\ \mathcal{G}_{1}U^{2}&\mathcal{G}_{2}U^{2}&\mathcal{G}_{3}U^{2}\\ \mathcal{G}_{1}U^{3}&\mathcal{G}_{2}U^{3}&\mathcal{G}_{3}U^{3}\end{matrix}\right]\right)^{\top}$ (84)
$=$ $\left[\begin{array}[]{ccc}{P}_{i1}\mathcal{G}_{1}&{P}_{i2}\mathcal{G}_{1}&{P}_{i3}\mathcal{G}_{1}\\ {P}_{i1}\mathcal{G}_{2}&{P}_{i2}\mathcal{G}_{2}&{P}_{i3}\mathcal{G}_{2}\\ {P}_{i1}\mathcal{G}_{3}&{P}_{i2}\mathcal{G}_{3}&{P}_{i3}\mathcal{G}_{3}\end{array}\right]\left[\begin{array}[]{c}U^{1}\\ U^{2}\\ U^{3}\end{array}\right]:=\mathcal{S}_{i}U.$
@@
$\Delta_{L}u=\mathcal{H}_{1}\left[\mathcal{H}_{1}+\mathcal{S}_{1}\right]U+\cdots+\mathcal{H}_{3}\left[\mathcal{H}_{3}+\mathcal{S}_{3}\right]U:=\bar{\Delta}_{L}U.$
@@
$\Delta_{L}u|_{S_{{\mathbf{x}_{0}}}}=\bar{\Delta}_{L}U|_{S_{{\mathbf{x}_{0}}}}=\sum_{\ell=1}^{3}\mathcal{H}_{\ell}\left(\mathcal{H}_{\ell}+\mathcal{S}_{\ell}\right)U|_{S_{{\mathbf{x}_{0}}}}\approx\sum_{\ell=1}^{3}\mathbf{H}_{\ell}\left(\mathbf{H}_{\ell}+\mathbf{S}_{\ell}\right)\mathbf{U}_{{\mathbf{x}_{0}}},$ (86)
@@
$\Delta_{L}u|_{S_{{\mathbf{x}_{0}}}}\approx\sum_{\ell=1}^{3}\mathbf{H}_{\ell}\left(\mathbf{H}_{\ell}+\mathbf{S}_{\ell}\right)\mathbf{U}_{{\mathbf{x}_{0}}}=\sum_{\ell=1}^{3}\mathbf{P}^{\otimes}\mathbf{H}_{\ell}\mathbf{P}^{\otimes}\left(\mathbf{H}_{\ell}+\mathbf{S}_{\ell}\right)\mathbf{P}^{\otimes}\mathbf{U}_{{\mathbf{x}_{0}}},$
@@
$\Delta_{L}u|_{S_{{\mathbf{x}_{0}}}}$$\approx$ $\sum_{\ell=1}^{3}\mathbf{T}^{\otimes}(\mathbf{T}^{\otimes\top}\mathbf{H}_{\ell}\mathbf{T}^{\otimes})[\mathbf{T}^{\otimes\top}\left(\mathbf{H}_{\ell}+\mathbf{S}_{\ell}\right)\mathbf{T}^{\otimes}](\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}})$ (87)
@@
where $\mathbf{R}_{\ell}:=\mathbf{T}^{\otimes\top}\mathbf{H}_{\ell}\mathbf{T}^{\otimes}=\mathbf{T}^{\otimes\top}\left(\mathbf{I}_{3}\mathbf{\otimes G}_{\ell}\right)\mathbf{T}^{\otimes}$ has been defined in (32), $\mathbf{M}_{\ell}:=\mathbf{T}^{\otimes\top}\mathbf{S}_{\ell}\mathbf{T}^{\otimes}$ is the reduction of $\mathbf{S}_{\ell}$, and $\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\left(\mathbf{R}_{\ell}+\mathbf{M}_{\ell}\right)\in\mathbb{R}^{2K\times 2K}$ is the L Laplacian matrix that maps the coordinate representation of $u|_{S_{{\mathbf{x}_{0}}}}$ to the coordinate representation of $(\Delta_{L}u)|_{S_{{\mathbf{x}_{0}}}}$ w.r.t. the global basis $\{\bm{t}_{j}(\mathbf{x}_{0,k})\}_{j=1,2}^{k=1,\ldots,K}$.
@@
Here, these weights satisfy $\Delta_{L}u({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{w}^{k}u({\mathbf{x}_{0,k}})$ at the base point $\mathbf{x}_{0}$.
@@ B.3 Hodge Laplacian
$\Delta_{H}u=\mathbf{P}\mathrm{tr}^{2}_{1}\left(\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}(V-V^{\top})\right)+\mathbf{P}\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}(\mathrm{tr}^{1}_{1}[\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}U]),$ (90)
where $V=\mathbf{P}\left(\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}U\right)\mathbf{P}$ is an extension of $v={\mathrm{grad}}_{g}u$ such that $V|_{M}=v$. We note that the first term in above (90) is only different from the extrinsic formulation for the L Laplacian in (77) by a negative sign. Thus, it only remains to compute the last term in (90),
$\mathbf{P}\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}(\mathrm{tr}^{1}_{1}[\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}U])=\mathbf{P}\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}(\sum_{k=1}^{3}\mathcal{G}_{k}U^{k})=\left[\begin{array}[]{c}\mathcal{G}_{1}\\ \mathcal{G}_{2}\\ \mathcal{G}_{3}\end{array}\right]\left[\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}\right]\left[\begin{array}[]{c}U^{1}\\ U^{2}\\ U^{3}\end{array}\right].$ (91)
@@
$\mathbf{P}\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}\left(\mathrm{tr}^{1}_{1}\left[\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}U\right]\right)|_{S_{\mathbf{x}_{0}}}$$\approx$ $\mathbf{T}^{\otimes}\left(\mathbf{T}^{\otimes\top}\left[\begin{array}[]{c}\mathbf{G}_{1}\\ \mathbf{G}_{2}\\ \mathbf{G}_{3}\end{array}\right]\right)\Big{(}\big{[}\mathbf{G}_{1},\mathbf{G}_{2},\mathbf{G}_{3}\big{]}\mathbf{T}^{\otimes}\Big{)}\Big{(}\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}\Big{)}$
@@
$\Delta_{H}u|_{S_{{\mathbf{x}_{0}}}}\approx\mathbf{T}^{\otimes}\left(\Big{[}\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\left(\mathbf{R}_{\ell}-\mathbf{M}_{\ell}\right)\Big{]}+\mathbf{J}\right)(\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}).$
@@
which are the $2$ rows of the matrix $\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\left(\mathbf{R}_{\ell}-\mathbf{M}_{\ell}\right)+\mathbf{J}$ corresponding to the base point $\mathbf{x}_{0}$ such that $\Delta_{H}u({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{w}^{k}u({\mathbf{x}_{0,k}})$.

### v2

@@ B.2 L Laplacian
We first review the extrinsic formulation for the L Laplacian (see e.g. [57, 24]). Let $\bm{v}={\mathrm{grad}}_{g}\bm{u}$ be the gradient of a vector field $\bm{u}$ and let $\bm{V}$ be a smooth extension of $\bm{v}$ onto an open subset in $\mathbb{R}^{3}$ such that $\bm{V}|_{M}=\bm{v}$. Based on the definition of the L Laplacian in (58), one can obtain the extrinsic formulation for the L Laplacian:
$\Delta_{L}\bm{u}$ $=$ $\mathrm{div}_{g}\left({\bm{v}+\bm{v}}^{\top}\right)=\mathbf{P}\mathrm{tr}^{2}_{1}\left(\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}(\bm{V}+\bm{V}^{\top})\right)=\mathbf{P}\left[\begin{array}[]{c}\sum_{k=1}^{3}\mathcal{G}_{k}\left(V^{1k}+V^{k1}\right)\\ \sum_{k=1}^{3}\mathcal{G}_{k}\left(V^{2k}+V^{k2}\right)\\ \sum_{k=1}^{3}\mathcal{G}_{k}\left(V^{3k}+V^{k3}\right)\end{array}\right]$ (92)
$=$ $\mathcal{H}_{1}\left[\bm{V}\mathbf{|}_{1}+\left(\bar{\bm{V}}_{1}\right)^{\top}\right]+\mathcal{H}_{2}\left[\bm{V}\mathbf{|}_{2}+\left(\bar{\bm{V}}_{2}\right)^{\top}\right]+\mathcal{H}_{3}\left[\bm{V}\mathbf{|}_{3}+\left(\bar{\bm{V}}_{3}\right)^{\top}\right],$ (93)
where $\bar{\bm{V}}_{i}=\left[V_{i1},V_{i2},V_{i3}\right]$ is the $i$th row of the $(2,0)$ tensor field $\bm{V}$ for $i=1,2,3$. Here, we have used the notations $\mathcal{G}_{j}$ in (34) and $\mathcal{H}_{j}$ in (37) for $j=1,2,3$, and we can further calculate $\bar{\bm{V}}^{i}$ as
$(\bar{\bm{V}}^{i})^{\top}$ $=$ $\left[V_{i1},V_{i2},V_{i3}\right]^{\top}=\left(\left[P_{i1},P_{i2},P_{i3}\right]\left[\begin{matrix}\mathcal{G}_{1}U^{1}&\mathcal{G}_{2}U^{1}&\mathcal{G}_{3}U^{1}\\ \mathcal{G}_{1}U^{2}&\mathcal{G}_{2}U^{2}&\mathcal{G}_{3}U^{2}\\ \mathcal{G}_{1}U^{3}&\mathcal{G}_{2}U^{3}&\mathcal{G}_{3}U^{3}\end{matrix}\right]\right)^{\top}$ (100)
$=$ $\left[\begin{array}[]{ccc}{P}_{i1}\mathcal{G}_{1}&{P}_{i2}\mathcal{G}_{1}&{P}_{i3}\mathcal{G}_{1}\\ {P}_{i1}\mathcal{G}_{2}&{P}_{i2}\mathcal{G}_{2}&{P}_{i3}\mathcal{G}_{2}\\ {P}_{i1}\mathcal{G}_{3}&{P}_{i2}\mathcal{G}_{3}&{P}_{i3}\mathcal{G}_{3}\end{array}\right]\left[\begin{array}[]{c}U^{1}\\ U^{2}\\ U^{3}\end{array}\right]:=\mathcal{S}_{i}\bm{U}.$
@@
$\Delta_{L}\bm{u}=\mathcal{H}_{1}\left[\mathcal{H}_{1}+\mathcal{S}_{1}\right]\bm{U}+\cdots+\mathcal{H}_{3}\left[\mathcal{H}_{3}+\mathcal{S}_{3}\right]\bm{U}:=\bar{\Delta}_{L}\bm{U}.$
@@
$(\Delta_{L}\bm{u})|_{S_{{\mathbf{x}_{0}}}}=(\bar{\Delta}_{L}\bm{U})|_{S_{{\mathbf{x}_{0}}}}=\sum_{\ell=1}^{3}(\mathcal{H}_{\ell}\left(\mathcal{H}_{\ell}+\mathcal{S}_{\ell}\right)\bm{U})|_{S_{{\mathbf{x}_{0}}}}\approx\sum_{\ell=1}^{3}\mathbf{H}_{\ell}\left(\mathbf{H}_{\ell}+\mathbf{S}_{\ell}\right)\mathbf{U}_{{\mathbf{x}_{0}}},$ (102)
@@
$(\Delta_{L}\bm{u})|_{S_{{\mathbf{x}_{0}}}}\approx\sum_{\ell=1}^{3}\mathbf{H}_{\ell}\left(\mathbf{H}_{\ell}+\mathbf{S}_{\ell}\right)\mathbf{U}_{{\mathbf{x}_{0}}}=\sum_{\ell=1}^{3}\mathbf{P}^{\otimes}\mathbf{H}_{\ell}\mathbf{P}^{\otimes}\left(\mathbf{H}_{\ell}+\mathbf{S}_{\ell}\right)\mathbf{P}^{\otimes}\mathbf{U}_{{\mathbf{x}_{0}}},$
@@
$\Delta_{L}\bm{u}|_{S_{{\mathbf{x}_{0}}}}$$\approx$ $\sum_{\ell=1}^{3}\mathbf{T}^{\otimes}(\mathbf{T}^{\otimes\top}\mathbf{H}_{\ell}\mathbf{T}^{\otimes})[\mathbf{T}^{\otimes\top}\left(\mathbf{H}_{\ell}+\mathbf{S}_{\ell}\right)\mathbf{T}^{\otimes}](\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}})$ (103)
@@
where $\mathbf{R}_{\ell}:=\mathbf{T}^{\otimes\top}\mathbf{H}_{\ell}\mathbf{T}^{\otimes}=\mathbf{T}^{\otimes\top}\left(\mathbf{I}_{3}\mathbf{\otimes G}_{\ell}\right)\mathbf{T}^{\otimes}$ has been defined in (44), $\mathbf{M}_{\ell}:=\mathbf{T}^{\otimes\top}\mathbf{S}_{\ell}\mathbf{T}^{\otimes}$ is the reduction of $\mathbf{S}_{\ell}$, and $\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\left(\mathbf{R}_{\ell}+\mathbf{M}_{\ell}\right)\in\mathbb{R}^{2K\times 2K}$ is the L Laplacian matrix that maps the coordinate representation of $\bm{u}|_{S_{{\mathbf{x}_{0}}}}$ to the coordinate representation of $(\Delta_{L}\bm{u})|_{S_{{\mathbf{x}_{0}}}}$ w.r.t. the global basis $\{\bm{t}_{j}(\mathbf{x}_{0,k})\}_{j=1,2}^{k=1,\ldots,K}$.
@@
Here, these weights satisfy $\Delta_{L}\bm{u}({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{w}^{k}\bm{u}|_{{\mathbf{x}_{0,k}}}$ at the base point $\mathbf{x}_{0}$.
@@ B.3 Hodge Laplacian
$\Delta_{H}\bm{u}=\mathbf{P}\mathrm{tr}^{2}_{1}\left(\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}(\bm{V}-\bm{V}^{\top})\right)+\mathbf{P}\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}(\mathrm{tr}^{1}_{1}[\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}\bm{U}]),$ (106)
where $\bm{V}=\mathbf{P}\left(\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}\bm{U}\right)\mathbf{P}$ is an extension of $\bm{v}={\mathrm{grad}}_{g}\bm{u}$ such that $\bm{V}|_{M}=\bm{v}$. We note that the first term in above (106) is only different from the extrinsic formulation for the L Laplacian in (93) by a negative sign. Thus, it only remains to compute the last term in (106),
$\mathbf{P}\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}(\mathrm{tr}^{1}_{1}[\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}\bm{U}])=\mathbf{P}\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}(\sum_{k=1}^{3}\mathcal{G}_{k}U^{k})=\left[\begin{array}[]{c}\mathcal{G}_{1}\\ \mathcal{G}_{2}\\ \mathcal{G}_{3}\end{array}\right]\left[\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}\right]\left[\begin{array}[]{c}U^{1}\\ U^{2}\\ U^{3}\end{array}\right].$ (107)
@@
$\mathbf{P}\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}\left(\mathrm{tr}^{1}_{1}\left[\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}\bm{U}\right]\right)|_{S_{\mathbf{x}_{0}}}$$\approx$ $\mathbf{T}^{\otimes}\left(\mathbf{T}^{\otimes\top}\left[\begin{array}[]{c}\mathbf{G}_{1}\\ \mathbf{G}_{2}\\ \mathbf{G}_{3}\end{array}\right]\right)\Big{(}\big{[}\mathbf{G}_{1},\mathbf{G}_{2},\mathbf{G}_{3}\big{]}\mathbf{T}^{\otimes}\Big{)}\Big{(}\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}\Big{)}$
@@
$(\Delta_{H}\bm{u})|_{S_{{\mathbf{x}_{0}}}}\approx\mathbf{T}^{\otimes}\left(\Big{[}\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\left(\mathbf{R}_{\ell}-\mathbf{M}_{\ell}\right)\Big{]}+\mathbf{J}\right)(\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}).$
@@
which are the $2$ rows of the matrix $\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\left(\mathbf{R}_{\ell}-\mathbf{M}_{\ell}\right)+\mathbf{J}$ corresponding to the base point $\mathbf{x}_{0}$ such that $\Delta_{H}\bm{u}({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{w}^{k}\bm{u}|_{{\mathbf{x}_{0,k}}}$.